OpenStudy (anonymous):

4x2 + 12x = 0 a) x = 3 b) x = 0, -3 c) x = -3 d) x = 0, 3

OpenStudy (anonymous):

b)

OpenStudy (anonymous):

use factoring...

OpenStudy (anonymous):

4x(x+3)=0

OpenStudy (anonymous):

x= 0, -3

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

b) x = 0, -3